OpenStudy (anonymous):

Write each # in the form a+bi 2+sqrt-3 sqrt-8+8

jimthompson5910 (jim_thompson5910):

What does \(\Large \sqrt{-3}\) simplify to?

OpenStudy (anonymous):

i(sqrt)3?

jimthompson5910 (jim_thompson5910):

good, so this means \[\Large 2+\sqrt{-3} = 2+i\sqrt{3}\]

jimthompson5910 (jim_thompson5910):

How about \(\Large \sqrt{-8}\) ?

OpenStudy (anonymous):

2isqrt2

jimthompson5910 (jim_thompson5910):

That means \[\Large \sqrt{-8} + 8 = 2i\sqrt{2} + 8\] \[\Large \sqrt{-8} + 8 = 8 + 2i\sqrt{2}\]

jimthompson5910 (jim_thompson5910):

The first answer \(\Large 2 + i\sqrt{3}\) can be written as \(\Large 2 + \sqrt{3} * i\) so it's in the form a+bi where a = 2 and \(\Large b = \sqrt{3}\)

jimthompson5910 (jim_thompson5910):

\(\Large 8 + 2i\sqrt{2}\) is the same as \(\Large 8 + 2\sqrt{2}*i\) so a = 8 and \(\Large b=2\sqrt{2}\)